For the last 8 years Paul has made deposits of $84.00 at the end of every month earning interest at 9% compounded monthly. If he leaves the accumulated balance for another 5 years at 5% compounded quarterly, what will the balance be in the account?

The balance will be $__?

(Round the final answer to the nearest cent as needed. Round all intermediate values to six decimal places as needed.)
Answer #1

Balance after 8 years of deposit=Future Value of annuity = A* ((1+rate)^n-1)/rate

= 84*((1+9%/12)^(8*12)-1)/(9%/12)

= 11747.917756

Final balance after 5 more years = Future value=Present value*(1+Rate)^number of periods

= 11747.917756*(1+5%/4)^(5*4)

= $15061.27